Circular No. 8631

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005ke IN NGC 1371 F. Patat and D. Baade, European Southern Observatory (ESO); L.

Wang, Lawrence Berkeley Laboratory; S. Taubenberger, Max-Planck-Insti-tut für Astrophysik, Garching; and J. C. Wheeler, University of Texas, report on observations of SN 2005ke (cf. IAUC 8630, CBET 289) on Nov. 16.12 UT with ESO’s Very Large Telescope (+ FORS1 in polarimetric mode). A fully reduced spectrum (range 340–855 nm, resolution 1.3 nm FWHM) shows 2005ke to be an underluminous type-Ia supernova before maximum, showing the characteristic 420-nm Ti II and 635-nm Si II lines. The spectral appearance is very similar to that of SN 1999by at 4 days before maximum light (Garnavich et al. 2004, Ap.J. 613, 1120). The expansion velocities deduced from the absorption minima of Si II, Ca II and O I are all around 11400 km/s — with the exception of Ca II H and K, from which a velocity of 13300 km/s is deduced. The synthetic colors derived from the spectrum are B − V = +0.33 and V − R = +0.24. A preliminary reduction of the spectropolarimetric data indicates a polarization degree of several percent in the Ca II infrared triplet. SN 2005ke was re-observed on Nov. 17.11 with the same instrumentation; the velocities have decreased by ∼ 400 km/s, the color became slightly redder (B − V = +0.45), and the polarization degree reduced by a factor of two. An earlier inspection of preliminarily reduced data led to the incorrect classification this object as a type-Ib/c supernova (cf. CBET 291).

COMET C/2005 W1 (SOHO) Further to IAUC 8626, K. Battams has reported measurements (initial

position given below) for a Marsden-group comet that was typically stellar but brightened significantly from mag nearly 9 when first detected in C3 images to ∼ 5.7 in C2 images. B. G. Marsden remarks on MPEC 2005-W07 (where he presents both parabolic and elliptical orbital elements for the currently visible comet) that this object is probably identical with C/2000 C4 (IAUC 7832, MPEC 2000-C52, MPC 44860), though it is conceivably identical with C/2000 C3 or C/2000 C7 (which both appeared around the same time — and, if they still exist, are also presumably about to return).

Comet  2005 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2005 W1 Nov. 16.579 15^{h}32.^{m}2 −20^{o}31′ C2/3 HS 2005-W07
2005 November 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green